# The effect of accretion on scalar superradiant instability

### v1

@@ II.2.1 The evolution without the scalar field
$k$ $=\frac{3a_{*0}}{\sqrt{\frac{12\sqrt{2}}{\sqrt{{3\beta^{4}}(1+a_{*0})^{-1}+3\beta^{2}+2}}+18-\gamma^{2}}+2\sqrt{6}-\gamma},$ (14a)
$\gamma$ $=\beta^{-1}\sqrt{9(1-a_{*0}^{2}+\beta^{4})+6\beta^{2}},$ (14b)
$\beta$ =($6]{(1+a_{*0})^{2}(1-a_{*0})}.$ (14c)
For the special case with $a_{*0}=0$, one obtains $k=1$ and the BH spin reaches its maximum value at $M=\sqrt{6}M_{0}$. For other values of $a_{*0}$, there is $10<k<$ 1. The BH with $a_{*0}<1$ evolves along the nontrivial solution, with both mass and spin increasing with time, until the spin reaches identity. Then the spin is fixed at 1 with mass increasing monotonically.
@@
$\frac{\dot{M}_{\mathrm{acc}}\eta}{1-\eta}=L_{\mathrm{Edd}}\equiv\frac{4\pi M\mu_{\mathrm{p}}}{\sigma_{\mathrm{T}}},$ (15)
where $\eta$ is the efficiency of photon radiation from the accretion disk, $\sigma_{\mathrm{T}}$ is the Thompson cross-section, and $\mu_{\mathrm{p}}$ is the proton mass. For a thin accretion disk, the efficiency has a simple form $\eta=1-E_{\mathrm{ms}}^{\dagger}$ Thorne:1974ve . For a Schwarzschild BH, $\eta$ is approximately $0.06$, while for an extreme Kerr BH with $a_{*}=1$, the value of $\eta$ can reach the value of $0.4$. Due to the small variation of $\eta$, it is usually assumed to be a constant for simplicity. According to Eq. (15), the BH mass then increases exponentially with a time scale
$\tau_{\mathrm{S}}\equiv\frac{\sigma_{\mathrm{T}}}{4\pi\mu_{\mathrm{p}}}\cdot\frac{\eta}{1-\eta}\approx 4.5\times 10^{8}\,\mathrm{yr}\cdot\frac{\eta}{1-\eta},$ (16)
which is also known as the Salpeter timescale Salpeter:1964kb .
@@
$\dot{M}_{\mathrm{acc}}=f_{\mathrm{Edd}}\frac{1-\eta}{\eta}L_{\mathrm{Edd}}\equiv M(t)/{\tau}_{\mathrm{acc}},$ (17)
@@
$\tau_{\mathrm{acc}}\equiv\tau_{\mathrm{S}}/f_{\mathrm{Edd}}=5\times 10^{7}\mathrm{yr}/{f}_{\mathrm{Edd}}.$ (18)
Inserting Eq. (17) into Eq. (11a), one obtains a BH mass increasing exponentially with time
$M(t)=M_{0}\exp(t/{\tau}_{\mathrm{acc}}).$ (19)
Figure 1: The evolution of BH spin $a_{*}$ as a function of the mass ratio $M/M_{0}$ without the scalar field. Two initial BH spins $a_{*0}=0$ and $a_{*0}=0.7$ are considered. The red dashed curves are solved from Eq. (20). The black curves further assumes the absence of the photon captured by the BH, which gives Eq. (13)
Next, we add the contribution of photon capture by the BH. It has been shown that the photon-capturing rate of the BH is greater for photons with negative angular momentum, i.e. angular momentum opposite to the BH spin, than for photons with positive angular momentum Godfrey:1970am . As a result, capturing photons emitted by the accretion disk slows down the increase of the BH spin. Mathematically, Eq. (12) should be modified as follows Thorne:1974ve :
$\frac{\mathrm{d}a_{*}}{\mathrm{d}\ln M}$ $=\frac{1}{M}\frac{J_{\mathrm{ms}}^{\dagger}+\dot{J}_{\mathrm{rad}}}{E_{\mathrm{ms}}^{\dagger}+\dot{M}_{\mathrm{rad}}}-2a_{*}.$ (20)
Using the expressions of $\dot{M}_{\mathrm{rad}}$ and $\dot{J}_{\mathrm{rad}}$ calculated in Refs. Page:1974he ; Thorne:1974ve , we solve the equation numerically and present the solutions in Fig. 1. The effect of photon-capturing is visible only when $a_{*}\gtrsim 0.97$. We also obtain a limiting value for the BH spin, ${a}_{\mathrm{*lim}}\approx 0.998189$, which is consistent with Ref. Thorne:1974ve . Since the photon capture’s contribution to BH spin is smaller than $0.3\%$, its effect is always ignored in the analytic approximation throughout this work.
II.2.2 The evolution without accretion
In the rest of this section, we study the scenario in the absence of accretion. Mathematically, the terms with $M_{\text{acc}}$ in Eqs. (11) are taken to be zero. In this part, we consider a BH with initial mass $M_{0}$ and spin $a_{*0}$. The initial mass of the $\{n,l,m\}$ mode is set as $M_{\text{s},0}^{(nlm)}$. The $a_{*0}$ is assumed to be larger than $a_{*\text{c}}^{(111)}$ such that the dominant $\{0,1,1\}$ mode and the subdominant $\{1,1,1\}$ mode grow due to superradiant instability. We focus on these two modes, with all other modes ignored. The evolution of this system can be separated into the spin-down phase, attractor phase of the subdominant mode, and attractor phase of the dominant mode. There has been extensive research on such isolated BH-condensate systems in the literature. Below, we summarize the findings in Ref. Guo:2022mpr .
@@ II.2.2 The evolution without accretion
$\begin{split}\frac{{M}_{\mathrm{s,1}}^{(011)}-{M}_{\mathrm{s,0}}^{(011)}}{M_{0}}&\approx M_{0}\mu\Big{[}a_{*0}-4M_{0}\mu+7.89a_{*0}(M_{0}\mu)^{2}\\ &\hskip 14.22636pt+(-16.55+6.45a_{*0}^{2})(M_{0}\mu)^{3}\Big{]},\end{split}$ (21)
$\frac{{M}_{\mathrm{s,1}}^{(111)}}{M_{0}}$ $\approx\frac{{M}_{\mathrm{s,0}}^{(111)}}{M_{0}}\left[\frac{{M}_{\mathrm{s,1}}^{(011)}}{{M}_{\mathrm{s,0}}^{(011)}}\right]^{0.35},$ (22)
where ${M}_{\mathrm{s,1}}^{(nlm)}\equiv{M}_{\mathrm{s}}^{(nlm)}(t_{1})$. The numbers in Eq. (21) are obtained by fitting the numerical result from solving the angular momentum and energy conservation equations
${a}_{\mathrm{*c}}^{(011)}M_{1}^{2}-a_{*0}M_{0}^{2}$ $=\frac{{M}_{\mathrm{s,1}}^{(011)}}{\omega_{011}(t_{1})}-\frac{{M}_{\mathrm{s,0}}^{(011)}}{\omega_{011}(t_{0})},$ (23a)
$M_{0}+{M}_{\mathrm{s,0}}^{(011)}$ $=M_{1}+{M}_{\mathrm{s,1}}^{(011)},$ (23b)
@@
With ${M}_{\mathrm{s,1}}^{(011)}$ obtained in Eq. (21), the $t_{1}$ could be estimated as
$t_{1}\approx\tau_{\text{SR}}^{(011)}\log\frac{M_{\mathrm{s,1}}^{(011)}}{M_{\mathrm{s,0}}^{(011)}},$ (24)
@@
$M\Gamma_{011}\approx\frac{1}{48}\left(a_{*}-4M\mu\right)(M\mu)^{9}.$ (25)
If satisfied with an order of magnitude estimate, one could further take the logarithm of the mass ratio in Eq. (24) to be roughly 10, and obtain $t_{1}$ in the SI units
$t_{1}\approx 5.09\,\mathrm{yr}\left(\frac{10M_{\odot}}{M_{0}}\right)^{8}\left(\frac{10^{-12}\,\mathrm{eV}}{\mu}\right)^{9}\left(\frac{1}{a_{*0}}\right).$ (26)
@@
${M}_{\mathrm{s}}^{(011)}\Gamma_{011}+{M}_{\mathrm{s}}^{(111)}\Gamma_{111}=0.$ (27)
@@
$\dot{E}_{\mathrm{SR}}^{(111)}=2{M}_{\mathrm{s}}^{(111)}\Gamma_{111}\approx-\frac{80}{19683}\frac{{M}_{\mathrm{s}}^{(111)}}{M_{1}}(M_{1}\mu)^{12},$ (28)
@@
$\dot{E}_{\mathrm{GW}}^{(111)}$ $\approx\frac{2\left(484+9\pi^{2}\right)}{32805}\frac{{M}_{\mathrm{s}}^{(111)}}{M_{1}}a_{*0}(M_{1}\mu)^{15},$ (29)
which is suppressed by $(M\mu)^{3}$ compared to $\dot{E}_{\text{SR}}$. Thus one could safely ignore the GW emission and use Eq. (28) to solve Eq. (11c). This observation applies to all $l>1$ modes. Then the scale of the $\{1,1,1\}$ mode lifetime could be derived as
${{\tau}_{\mathrm{life}}^{(111)}}$ $=\frac{19683}{80}M_{1}(M_{1}\mu)^{-12}.$ (30)
@@
The attractor phase of the dominant mode starts from $t_{2}$. In this phase, the BH has constant mass $M_{1}$ and spin $a_{*\text{c}}^{(011)}$. The GW emission dominates the evolution of the $\{0,1,1\}$ mode. One could then solve Eq. (11c) directly and obtain
${M}_{\mathrm{s}}^{(011)}(t)=\frac{{M}_{\mathrm{s,2}}^{(011)}}{1+(t-t_{2})/{\tau}_{\mathrm{GW}}^{(011)}},$ (31)
@@
$\begin{split}{{\tau}_{\mathrm{GW}}^{(011)}}&=40.22\cdot\frac{M_{1}^{2}}{{M}_{\mathrm{s,2}}^{(011)}}(M_{1}\mu)^{-14}.\end{split}$ (32)
@@
$\tau_{\mathrm{GW}}^{(011)}\approx 4.85\times 10^{6}\mathrm{yr}\left(\frac{10M_{\odot}}{M_{0}}\right)^{14}\left(\frac{10^{-12}\mathrm{eV}}{\mu}\right)^{15}\left(\frac{1}{a_{*0}}\right).$ (33)
@@
In a realistic case, the $l=m=2$ modes become important when the $\{0,1,1\}$ mode is almost depleted via GW emission. The latter evolution is very similar to the process with $l=m=1$ modes described above. The analytic estimates are still valid for the $l=m=2$ modes, with the initial BH spin and mass now replaced by ${a}_{\mathrm{*c}}^{(011)}$ and $M_{1}$, respectively. More details can be found in Ref. Guo:2022mpr .
III Effect of accretion on single mode evolution
@@
In this section, we study the effect of accretion on the mode $\{0,1,1\}$, which has the largest superradiance rate. Other modes are ignored at the moment. The important time scales are analyzed in Sec. III.1. If the superradiance could happen initially without accretion, the evolution further depends on the relative size of these time scales, which are explained below in Sec. III.2-III.4. If the superradiance is forbidden initially, the evolution is a little different, which is studied in Sec. III.5. Finally, we summarize this section and point out the universal behavior of the BH evolution in Sec. III.6.
III.1 Time Scales
@@
If superradiance could happen without accretion, there are two widely separated scales: the superradiance time scale $\tau_{\text{SR}}$ and the GW emission time scale $\tau_{\text{GW}}$, with the latter much larger than the former. With accretion, there is a new time scale $\tau_{\text{acc}}$. Another time scale can be found by requiring the critical BH spin in Eq. (10) to be less than the limiting value 0.998, which gives a critical value of the BH mass
$M_{\text{de}}^{(nlm)}\approx\frac{0.47m}{\mu},$ (34)
@@
where the subscript stands for “decaying”. Once the BH mass surpasses this critical value, the superradiant mode turns to a decaying mode. In this section, we consider the case with $M_{0}<M_{\text{de}}^{(011)}$. We refer to the time cost as the decaying time scale $\tau_{\text{de}}$. Using Eq. (19), one could estimate this decaying time scale as
$\tau_{\text{de}}=\tau_{\text{acc}}\log\frac{0.47m}{\alpha_{0}},$ (35)
@@
where $\alpha_{0}=M_{0}\mu$. For the phenomenologically important cases, the value of $\alpha_{0}$ is roughly $0.01$ and the $\tau_{\text{de}}$ is in the same order as $\tau_{\text{acc}}$. Thus it is not considered as a separate time scale below.
@@
The value of $\tau_{\text{acc}}$ can be adjusted by changing the accretion rate $f_{\text{Edd}}$ as in Eq. (18). There are five scenarios with this new scale, i,e, ${\tau}_{\mathrm{acc}}\ll{\tau}_{\mathrm{SR}}^{(011)}$, ${\tau}_{\mathrm{acc}}\sim{\tau}_{\mathrm{SR}}^{(011)}$, ${\tau}_{\mathrm{SR}}^{(011)}\ll{\tau}_{\mathrm{acc}}\ll{\tau}_{\mathrm{GW}}^{(011)}$, ${\tau}_{\mathrm{acc}}\sim{\tau}_{\mathrm{GW}}^{(011)}$ and ${\tau}_{\mathrm{acc}}\gg{\tau}_{\mathrm{GW}}^{(011)}$. In Fig. 2, we numerically solve Eqs. (11) for these five scenarios. The initial BH mass and spin are chosen as $M_{0}=10^{3}M_{\odot}$ and $a_{*0}=0.7$, respectively. The initial mass coupling is set as $M_{0}\mu=0.01$, which then gives the scalar mass $\mu\approx 1.34\times 10^{-15}\,\mathrm{eV}$. The superradiance and the GW emission time scales without accretion are ${\tau}_{\mathrm{SR}}^{(011)}\approx 6.01\times 10^{9}\,\mathrm{yr}$ and ${\tau}_{\mathrm{GW}}^{(011)}\approx 1.03\times 10^{22}\,\mathrm{yr}$. Their values are plotted in Fig. 2 as vertical lines. The case without accretion ($f_{\text{Edd}}=0$) is plotted in dashed curves for comparison. One can see that the BH-condensate system evolution is dramatically affected by a large accretion rate. Specifically, both the superradiance and the GW emission proceed much faster than the case with $f_{\text{Edd}}=0$. By observing the curves in Fig. 2, we further categorize the above five scenarios into three groups, which are studied in the next three subsections respectively.
@@
On the other hand, if the superradiance could not happen initially, neither the superradiance nor the GW emission time scales exist. The BH evolution is controlled by the accretion until the extraction of the angular momentum is faster than the inwards flow due to the accretion. This case is studied in Sec. III.5.
@@
III.2 $a_{*0}>a_{*\text{c}}^{(001)}$ and ${\tau}_{\mathrm{acc}}\ll{\tau}_{\mathrm{SR}}^{(011)}$
The case of ${\tau}_{\mathrm{acc}}\ll{\tau}_{\mathrm{SR}}^{(011)}$ corresponds to the curves with ${f}_{\mathrm{Edd}}=1$ in Fig. 2. They are re-plotted with more details in Fig. 3. The evolution was split into four distinct phases in Ref. Guo:2022mpr . Below we briefly review these phases: [^2] Here we rename the phases of those in Ref. Guo:2022mpr .
Figure 3: The time evolution of the scalar condensate mass (panel a), the GW emission (panel b), the BH spin (panel c), and the BH mass (panel d). The black and orange dotted curves are the evolution with and without the GW emission, respectively. The green curve in panel (c) is calculated with Eq. (10). The blue line in panel (d) is the exponential increase of the BH mass given in Eq. (19).
  • Spin-up phase. This phase corresponds to the interval $0<t<t_{0}^{\prime}$, where $t_{0}^{\prime}$ represents the time at which the BH spin reaches its first local maximum. During this phase, the BH evolves following the process described in Sec. II.2.1, while the scalar condensate is too small in mass to be important.
  • Spin-down phase. This phase spans $t_{0}^{\prime}<t<t_{1}$, where $t_{1}$ marks the time at which the BH spin reaches its minimum. As the BH mass and the condensate mass increase with time, the angular momentum extraction rate grows faster than exponentially. At some point, the extraction surpasses the accretion, resulting in a decrease of $a_{*}$ until it approaches the $\{0,1,1\}$ Regge trajectory determined by Eq. (10). This Regge trajectory is plotted as the green curve in Fig. 3(c).
  • Attractor phase. This phase extends from $t_{1}$ to $t_{3}$, where $t_{3}$ denotes the moment when the BH mass reaches $M_{\text{de}}^{(011)}$ given in Eq. (34). Throughout this phase, the BH evolves along the Regge trajectory of the $\{0,1,1\}$ mode, which acts as an attractor mathematically. Concurrently, as the condensate accumulates mass, its GW emission intensifies, and ${M}_{\mathrm{s}}^{(011)}/M$ reaches its maximum at $t_{2}$. Note that it does not coincide with the peak of the GW emission flux in panel (b), since the latter depends on both the mass ratio and the mass coupling, illustrated in Eq. (82).

### v2

@@ 2.2.1 The evolution without the scalar field
$k$ $=\frac{3a_{*0}}{\sqrt{\frac{12\sqrt{2}}{\sqrt{{3\beta^{4}}(1+a_{*0})^{-1}+3\beta^{2}+2}}+18-\gamma^{2}}+2\sqrt{6}-\gamma},$ (2.14a)
$\gamma$ $=\beta^{-1}\sqrt{9(1-a_{*0}^{2}+\beta^{4})+6\beta^{2}},$ (2.14b)
$\beta$ $=\sqrt[6]{(1+a_{*0})^{2}(1-a_{*0})}.$ (2.14c)
For the special case with $a_{*0}=0$, one obtains $k=1$ and the BH spin reaches its maximum value at $M=\sqrt{6}M_{0}$. For other values of $a_{*0}$, there is $0<k<1$. The BH with $a_{*0}<1$ evolves along the nontrivial solution, with both mass and spin increasing with time, until the spin reaches 1. Then the spin is fixed at 1 with mass increasing monotonically.
@@
$\frac{\dot{M}_{\mathrm{acc}}\eta}{1-\eta}=L_{\mathrm{Edd}}\equiv\frac{4\pi M\mu_{\mathrm{p}}}{\sigma_{\mathrm{T}}},$ (2.15)
where $\eta$ is the efficiency of photon radiation from the accretion disk, $\sigma_{\mathrm{T}}$ is the Thompson cross-section, and $\mu_{\mathrm{p}}$ is the proton mass. For a thin accretion disk, the efficiency has a simple form $\eta=1-E_{\mathrm{ms}}^{\dagger}$ [34]. For a Schwarzschild BH, $\eta$ is approximately $0.06$, while for an extreme Kerr BH with $a_{*}=1$, the value of $\eta$ can reach the value of $0.4$. Due to the small variation of $\eta$, it is usually assumed to be a constant for simplicity. According to Eq. (2.15), the BH mass then increases exponentially with a time scale
$\tau_{\mathrm{S}}\equiv\frac{\sigma_{\mathrm{T}}}{4\pi\mu_{\mathrm{p}}}\cdot\frac{\eta}{1-\eta}\approx 4.5\times 10^{8}\,\mathrm{yr}\cdot\frac{\eta}{1-\eta},$ (2.16)
which is also known as the Salpeter timescale [36].
@@
$\dot{M}_{\mathrm{acc}}=f_{\mathrm{Edd}}\frac{1-\eta}{\eta}L_{\mathrm{Edd}}\equiv M(t)/{\tau}_{\mathrm{acc}},$ (2.17)
@@
$\tau_{\mathrm{acc}}\equiv\tau_{\mathrm{S}}/f_{\mathrm{Edd}}=5\times 10^{7}\mathrm{yr}/{f}_{\mathrm{Edd}}.$ (2.18)
Inserting Eq. (2.17) into Eq. (2.11a), one obtains a BH mass increasing exponentially with time
$M(t)=M_{0}\exp(t/{\tau}_{\mathrm{acc}}).$ (2.19)
Figure 1: The evolution of BH spin $a_{*}$ as a function of the mass ratio $M/M_{0}$ without the scalar field. Two initial BH spins $a_{*0}=0$ and $a_{*0}=0.7$ are considered. The red dashed curves are solved from Eq. (2.20). The black curves further assumes the absence of the photon captured by the BH, which gives Eq. (2.13)
Next, we add the contribution of photon capture by the BH. It has been shown that the photon-capturing rate of the BH is greater for photons with negative angular momentum, i.e. angular momentum opposite to the BH spin, than for photons with positive angular momentum [37]. As a result, capturing photons emitted by the accretion disk slows down the increase of the BH spin. Mathematically, Eq. (2.12) should be modified as follows [34]:
$\frac{\mathrm{d}a_{*}}{\mathrm{d}\ln M}$ $=\frac{1}{M}\frac{J_{\mathrm{ms}}^{\dagger}+\dot{J}_{\mathrm{rad}}}{E_{\mathrm{ms}}^{\dagger}+\dot{M}_{\mathrm{rad}}}-2a_{*}.$ (2.20)
Using the expressions of $\dot{M}_{\mathrm{rad}}$ and $\dot{J}_{\mathrm{rad}}$ calculated in Refs. [35, 34], we solve the equation numerically and present the solutions in Fig. 1. The effect of photon-capturing is visible only when $a_{*}\gtrsim 0.97$. We also obtain a limiting value for the BH spin, ${a}_{\mathrm{*lim}}\approx 0.998189$, which is consistent with Ref. [34]. Since the photon capture’s contribution to BH spin is smaller than $0.3\%$, its effect is always ignored in the analytic approximation throughout this work.
2.2.2 The evolution without accretion
In the rest of this section, we study the scenario in the absence of accretion. Mathematically, the terms with $M_{\text{acc}}$ in Eqs. (2.11) are taken to be zero. In this part, we consider a BH with initial mass $M_{0}$ and spin $a_{*0}$. The initial mass of the $\{n,l,m\}$ mode is set as $M_{\text{s},0}^{(nlm)}$. The $a_{*0}$ is assumed to be larger than $a_{*\text{c}}^{(111)}$ such that the dominant $\{0,1,1\}$ mode and the subdominant $\{1,1,1\}$ mode grow due to superradiant instability. We focus on these two modes, with all other modes ignored. The evolution of this system can be separated into the spin-down phase, attractor phase of the subdominant mode, and attractor phase of the dominant mode. There has been extensive research on such isolated BH-condensate systems in the literature. Below, we summarize the findings in Ref. [16].
@@ 2.2.2 The evolution without accretion
$\frac{{M}_{\mathrm{s,1}}^{(011)}-{M}_{\mathrm{s,0}}^{(011)}}{M_{0}}$ $\approx M_{0}\mu\left[a_{*0}-4M_{0}\mu+7.89a_{*0}(M_{0}\mu)^{2}+(-16.55+6.45a_{*0}^{2})(M_{0}\mu)^{3}\right],$ (2.21)
$\frac{{M}_{\mathrm{s,1}}^{(111)}}{M_{0}}$ $\approx\frac{{M}_{\mathrm{s,0}}^{(111)}}{M_{0}}\left[\frac{{M}_{\mathrm{s,1}}^{(011)}}{{M}_{\mathrm{s,0}}^{(011)}}\right]^{0.35},$ (2.22)
where ${M}_{\mathrm{s,1}}^{(nlm)}\equiv{M}_{\mathrm{s}}^{(nlm)}(t_{1})$. The numbers in Eq. (2.21) are obtained by fitting the numerical result from solving the angular momentum and energy conservation equations
${a}_{\mathrm{*c}}^{(011)}M_{1}^{2}-a_{*0}M_{0}^{2}$ $=\frac{{M}_{\mathrm{s,1}}^{(011)}}{\omega_{011}(t_{1})}-\frac{{M}_{\mathrm{s,0}}^{(011)}}{\omega_{011}(t_{0})},$ (2.23a)
$M_{0}+{M}_{\mathrm{s,0}}^{(011)}$ $=M_{1}+{M}_{\mathrm{s,1}}^{(011)},$ (2.23b)
@@
With ${M}_{\mathrm{s,1}}^{(011)}$ obtained in Eq. (2.21), the $t_{1}$ could be estimated as
$t_{1}\approx\tau_{\text{SR}}^{(011)}\log\frac{M_{\mathrm{s,1}}^{(011)}}{M_{\mathrm{s,0}}^{(011)}},$ (2.24)
@@
$M\Gamma_{011}\approx\frac{1}{48}\left(a_{*}-4M\mu\right)(M\mu)^{9}.$ (2.25)
If satisfied with an order of magnitude estimate, one could further take the logarithm of the mass ratio in Eq. (2.24) to be roughly 10, and obtain $t_{1}$ in the SI units
$t_{1}\approx 5.09\,\mathrm{yr}\left(\frac{10M_{\odot}}{M_{0}}\right)^{8}\left(\frac{10^{-12}\,\mathrm{eV}}{\mu}\right)^{9}\left(\frac{1}{a_{*0}}\right).$ (2.26)
@@
${M}_{\mathrm{s}}^{(011)}\Gamma_{011}+{M}_{\mathrm{s}}^{(111)}\Gamma_{111}=0.$ (2.27)
@@
$\dot{E}_{\mathrm{SR}}^{(111)}=2{M}_{\mathrm{s}}^{(111)}\Gamma_{111}\approx-\frac{80}{19683}\frac{{M}_{\mathrm{s}}^{(111)}}{M_{1}}(M_{1}\mu)^{12},$ (2.28)
@@
$\dot{E}_{\mathrm{GW}}^{(111)}$ $\approx\frac{2\left(484+9\pi^{2}\right)}{32805}\frac{{M}_{\mathrm{s}}^{(111)}}{M_{1}}a_{*0}(M_{1}\mu)^{15},$ (2.29)
which is suppressed by $(M\mu)^{3}$ compared to $\dot{E}_{\text{SR}}$. Thus one could safely ignore the GW emission and use Eq. (2.28) to solve Eq. (2.11c). This observation applies to all $l>1$ modes. Then the scale of the $\{1,1,1\}$ mode lifetime could be derived as
${{\tau}_{\mathrm{life}}^{(111)}}$ $=\frac{19683}{80}M_{1}(M_{1}\mu)^{-12}.$ (2.30)
@@
The attractor phase of the dominant mode starts from $t_{2}$. In this phase, the BH has constant mass $M_{1}$ and spin $a_{*\text{c}}^{(011)}$. The GW emission dominates the evolution of the $\{0,1,1\}$ mode. One could then solve Eq. (2.11c) directly and obtain
${M}_{\mathrm{s}}^{(011)}(t)=\frac{{M}_{\mathrm{s,2}}^{(011)}}{1+(t-t_{2})/{\tau}_{\mathrm{GW}}^{(011)}},$ (2.31)
@@
${{\tau}_{\mathrm{GW}}^{(011)}}$ $=40.22\cdot\frac{M_{1}^{2}}{{M}_{\mathrm{s,2}}^{(011)}}(M_{1}\mu)^{-14}.$ (2.32)
@@
$\tau_{\mathrm{GW}}^{(011)}\approx 4.85\times 10^{6}\mathrm{yr}\left(\frac{10M_{\odot}}{M_{0}}\right)^{14}\left(\frac{10^{-12}\mathrm{eV}}{\mu}\right)^{15}\left(\frac{1}{a_{*0}}\right).$ (2.33)
@@
In a realistic case, the $l=m=2$ modes become important when the $\{0,1,1\}$ mode is almost depleted via GW emission. The latter evolution is very similar to the process with $l=m=1$ modes described above. The analytic estimates are still valid for the $l=m=2$ modes, with the initial BH spin and mass now replaced by ${a}_{\mathrm{*c}}^{(011)}$ and $M_{1}$, respectively. More details can be found in Ref. [16].
3 Effect of accretion on single mode evolution
@@
In this section, we study the effect of accretion on the mode $\{0,1,1\}$, which has the largest superradiance rate. Other modes are ignored at the moment. The important time scales are analyzed in Sec. 3.1. If the superradiance could happen initially without accretion, the evolution further depends on the relative size of these time scales, which are explained below in Sec. 3.2-3.4. If the superradiance is forbidden initially, the evolution is a little different, which is studied in Sec. 3.5. Finally, we summarize this section and point out the universal behavior of the BH evolution in Sec. 3.6.
3.1 Time Scales
@@
If superradiance could happen without accretion, there are two widely separated scales: the superradiance time scale $\tau_{\text{SR}}$ and the GW emission time scale $\tau_{\text{GW}}$, with the latter much larger than the former. With accretion, there is a new time scale $\tau_{\text{acc}}$. Another time scale can be found by requiring the critical BH spin in Eq. (2.10) to be less than the limiting value 0.998, which gives a critical value of the BH mass
$M_{\text{de}}^{(nlm)}\approx\frac{0.47m}{\mu},$ (3.1)
@@
where the subscript stands for “decaying”. Once the BH mass surpasses this critical value, the superradiant mode turns to a decaying mode. In this section, we consider the case with $M_{0}<M_{\text{de}}^{(011)}$. We refer to the time cost as the decaying time scale $\tau_{\text{de}}$. Using Eq. (2.19), one could estimate this decaying time scale as
$\tau_{\text{de}}=\tau_{\text{acc}}\log\frac{0.47m}{\alpha_{0}},$ (3.2)
@@
where $\alpha_{0}=M_{0}\mu$. For phenomenologically important cases, the value of $\alpha_{0}$ is roughly $0.01$, making the logarithm of order $\mathcal{O}(1)$. Thus, $\tau_{\text{de}}$ is of the same order as $\tau_{\text{acc}}$ and is not a separate time scale.
@@
The value of $\tau_{\text{acc}}$ can be adjusted by changing the accretion rate $f_{\text{Edd}}$ as in Eq. (2.18). There are five scenarios with this new scale, i,e, ${\tau}_{\mathrm{acc}}\ll{\tau}_{\mathrm{SR}}^{(011)}$, ${\tau}_{\mathrm{acc}}\sim{\tau}_{\mathrm{SR}}^{(011)}$, ${\tau}_{\mathrm{SR}}^{(011)}\ll{\tau}_{\mathrm{acc}}\ll{\tau}_{\mathrm{GW}}^{(011)}$, ${\tau}_{\mathrm{acc}}\sim{\tau}_{\mathrm{GW}}^{(011)}$ and ${\tau}_{\mathrm{acc}}\gg{\tau}_{\mathrm{GW}}^{(011)}$. In Fig. 2, we numerically solve Eqs. (2.11) for these five scenarios. The initial BH mass and spin are chosen as $M_{0}=10^{3}M_{\odot}$ and $a_{*0}=0.7$, respectively. The initial mass coupling is set as $M_{0}\mu=0.01$, which then gives the scalar mass $\mu\approx 1.34\times 10^{-15}\,\mathrm{eV}$. The superradiance and the GW emission time scales without accretion are ${\tau}_{\mathrm{SR}}^{(011)}\approx 6.01\times 10^{9}\,\mathrm{yr}$ and ${\tau}_{\mathrm{GW}}^{(011)}\approx 1.03\times 10^{22}\,\mathrm{yr}$. Their values are plotted in Fig. 2 as vertical lines. The case without accretion ($f_{\text{Edd}}=0$) is plotted in dashed curves for comparison. One can see that the BH-condensate system evolution is dramatically affected by a large accretion rate. Specifically, both the superradiance and the GW emission proceed much faster than the case with $f_{\text{Edd}}=0$. By observing the curves in Fig. 2, we further categorize the above five scenarios into three groups, which are studied in the next three subsections respectively.
@@
On the other hand, if the superradiance could not happen initially, neither the superradiance nor the GW emission time scales exist. The BH evolution is controlled by the accretion until the extraction of the angular momentum is faster than the inwards flow due to the accretion. This case is studied in Sec. 3.5.
@@
3.2 $a_{*0}>a_{*\text{c}}^{(001)}$ and ${\tau}_{\mathrm{acc}}\ll{\tau}_{\mathrm{SR}}^{(011)}$
The case of ${\tau}_{\mathrm{acc}}\ll{\tau}_{\mathrm{SR}}^{(011)}$ corresponds to the curves with ${f}_{\mathrm{Edd}}=1$ in Fig. 2. They are re-plotted with more details in Fig. 3. The evolution was split into four distinct phases in Ref. [16]. Below we briefly review these phases: [^2] Here we rename the phases of those in Ref. [16].
Figure 3: The time evolution of the scalar condensate mass (panel a), the GW emission (panel b), the BH spin (panel c), and the BH mass (panel d). The black and orange dotted curves are the evolution with and without the GW emission, respectively. The green curve in panel (c) is calculated with Eq. (2.10). The blue line in panel (d) is the exponential increase of the BH mass given in Eq. (2.19).
  • Spin-up phase. This phase corresponds to the interval $0<t<t_{0}^{\prime}$, where $t_{0}^{\prime}$ represents the time at which the BH spin reaches its first local maximum. During this phase, the BH evolves following the process described in Sec. 2.2.1, while the scalar condensate is too small in mass to be important.
  • Spin-down phase. This phase spans $t_{0}^{\prime}<t<t_{1}$, where $t_{1}$ marks the time at which the BH spin reaches its minimum. As the BH mass and the condensate mass increase with time, the angular momentum extraction rate grows faster than exponentially. At some point, the extraction surpasses the accretion, resulting in a decrease of $a_{*}$ until it approaches the $\{0,1,1\}$ Regge trajectory determined by Eq. (2.10). This Regge trajectory is plotted as the green curve in Fig. 3(c).
  • Attractor phase. This phase extends from $t_{1}$ to $t_{3}$, where $t_{3}$ denotes the moment when the BH mass reaches $M_{\text{de}}^{(011)}$ given in Eq. (3.1). Throughout this phase, the BH evolves along the Regge trajectory of the $\{0,1,1\}$ mode, which acts as an attractor mathematically. Concurrently, as the condensate accumulates mass, its GW emission intensifies, and ${M}_{\mathrm{s}}^{(011)}/M$ reaches its maximum at $t_{2}$. Note that it does not coincide with the peak of the GW emission flux in panel (b), since the latter depends on both the mass ratio and the mass coupling, illustrated in Eq. (A.5).